Determine an exponential function going through the points (-1,2)(3/2,1/2)

Answer by stanbon(75887)

Determine an exponential function going through the points (-1,2)(3/2,1/2)
-----
Form::
y = ab^x
--------
Equations:
2 = ab^(-1)
1/2 = ab^(3/2)
--------------------
Divide the bottom equation by the top equation:
1/4 = b^(5/2)
Then b = (1/4)^(2/5) = 0.5743
----------------------------
Now, solve for "a"::
2 = a*0.5743^(-1)
a = 2*0.5743
a = 1.1487
-----------------------
Equation::
y = 1.1487*(0.5743)^x
------------------------------
Cheers,
Stan H.
--------------